# On the Theoretical Foundation of Sparse Dictionary Learning in Mechanistic Interpretability

### v2

@@ 7 Conclusion
We construct a point $(W_{d},W_{e})$ exhibiting the polysemanticity pattern and verify it satisfies all three conditions.
Step 1: Construction of encoder $W_{e}$. Given a realizable polysemanticity pattern $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$, by Definition LABEL:def:realizable-pattern, there exists an encoder $W_{e}=[(w_{e}^{1})^{\top},\ldots,(w_{e}^{n_{q}})^{\top}]^{\top}$ such that for all $i\in[n_{q}]$:
$\mathcal{F}_{i}=\left\{d\in[n]:\left(\sigma(W_{e}w_{p}^{d})\right)_{i}>0\right\}.$ (43)
Fix this $W_{e}$ for the remainder of the proof. Define the activation vectors:
$z_{d}:=\sigma(W_{e}w_{p}^{d})\in\mathbb{R}^{n_{q}},\quad d=1,\ldots,n.$ (44)
By construction, $(z_{d})_{i}>0$ if and only if $d\in\mathcal{F}_{i}$.
Partition the neurons into active and dead sets:
$\mathcal{A}=\{i\in[n_{q}]:\mathcal{F}_{i}\neq\emptyset\},\quad\mathcal{D}=\{i\in[n_{q}]:\mathcal{F}_{i}=\emptyset\}.$ (45)
Step 2: Gradient analysis with respect to $W_{d}$. The gradient of the loss with respect to the $i$-th decoder column $w_{d}^{i}\in\mathbb{R}^{n_{r}}$ is:
$\nabla_{w_{d}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}=-2\sum_{d=1}^{n}M_{d}(w_{r}^{d}-W_{d}z_{d})(z_{d})_{i}.$ (46)
Case 1: Dead neurons ($i\in\mathcal{D}$). For $i\in\mathcal{D}$, we have $\mathcal{F}_{i}=\emptyset$, which implies $(z_{d})_{i}=0$ for all $d\in[n]$. Therefore:
$\nabla_{w_{d}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}=0\quad\forall w_{d}^{i}.$ (47)
The decoder weights for dead neurons are arbitrary. We set $w_{d}^{i}=0$ for all $i\in\mathcal{D}$.
Case 2: Active neurons ($i\in\mathcal{A}$). For $i\in\mathcal{A}$, we have $\mathcal{F}_{i}\neq\emptyset$. The gradient vanishes when:
$\sum_{d\in\mathcal{F}_{i}}M_{d}w_{r}^{d}(z_{d})_{i}=\sum_{d\in\mathcal{F}_{i}}M_{d}(W_{d}z_{d})(z_{d})_{i}.$ (48)
Expanding the right side and using the fact that $(z_{d})_{j}=0$ for $j\notin\mathcal{F}_{j}$ or $d\notin\mathcal{F}_{j}$:
$\sum_{d\in\mathcal{F}_{i}}M_{d}w_{r}^{d}(z_{d})_{i}=\sum_{j=1}^{n_{q}}w_{d}^{j}\underbrace{\sum_{d\in\mathcal{F}_{i}\cap\mathcal{F}_{j}}M_{d}(z_{d})_{i}(z_{d})_{j}}_{=:A_{ij}}.$ (49)
Define the Gram matrix $A\in\mathbb{R}^{n_{q}\times n_{q}}$:
$A_{ij}=\sum_{d\in\mathcal{F}_{i}\cap\mathcal{F}_{j}}M_{d}(z_{d})_{i}(z_{d})_{j}.$ (50)
And the target matrix $B\in\mathbb{R}^{n_{r}\times n_{q}}$:
$B_{:,i}=\sum_{d\in\mathcal{F}_{i}}M_{d}w_{r}^{d}(z_{d})_{i}.$ (51)
The stationarity condition becomes:
$B=W_{d}A.$ (52)
Step 3: Properties of the Gram matrix $A$.
Claim 3.1: Block structure. The matrix $A$ has a block-diagonal structure with respect to the partition $(\mathcal{A},\mathcal{D})$:
$A=\begin{bmatrix}A_{\mathcal{A}\mathcal{A}}&0\\ 0&0\end{bmatrix}$, (53)
where $A_{\mathcal{A}\mathcal{A}}$ corresponds to rows and columns indexed by $\mathcal{A}$.
Proof of Claim 3.1: For $i\in\mathcal{D}$, we have $\mathcal{F}_{i}=\emptyset$, so $\mathcal{F}_{i}\cap\mathcal{F}_{j}=\emptyset$ for all $j$, giving $A_{ij}=0$. Similarly, $A_{ji}=0$. For $i,j\in\mathcal{D}$, $A_{ij}=0$. $\square$
Claim 3.2: Positive definiteness. The submatrix $A_{\mathcal{A}\mathcal{A}}$ is positive definite.
Proof of Claim 3.2: For any nonzero $v\in\mathbb{R}^{|\mathcal{A}|}$, consider:
$v^{\top}A_{\mathcal{A}\mathcal{A}}v$ $=\sum_{i,j\in\mathcal{A}}v_{i}v_{j}A_{ij}$ (54)
$=\sum_{i,j\in\mathcal{A}}v_{i}v_{j}\sum_{d\in\mathcal{F}_{i}\cap\mathcal{F}_{j}}M_{d}(z_{d})_{i}(z_{d})_{j}$ (55)
$=\sum_{d=1}^{n}M_{d}\left(\sum_{i\in\mathcal{A}:d\in\mathcal{F}_{i}}v_{i}(z_{d})_{i}\right)^{2}.$ (56)
Since $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ forms a partition of $[n]$, for each $d\in[n]$, there exists at least one $i\in\mathcal{A}$ with $d\in\mathcal{F}_{i}$ and $(z_{d})_{i}>0$. For any nonzero $v$, at least one term in the sum is positive, giving $v^{\top}A_{\mathcal{A}\mathcal{A}}v>0$. $\square$
Step 4: Optimal decoder solution. From the normal equations (52), the optimal decoder for active neurons is:
$W_{d}^{*}[:,\mathcal{A}]=B[:,\mathcal{A}](A_{\mathcal{A}\mathcal{A}})^{-1},$ (57)
where $(A_{\mathcal{A}\mathcal{A}})^{-1}$ exists by Claim 3.2. For dead neurons, we set:
$W_{d}^{*}[:,\mathcal{D}]=0.$ (58)
This gives $\nabla_{W_{d}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})=0$, proving condition (2).
Step 5: Hessian analysis. The Hessian with respect to $W_{d}$ has Kronecker structure:
$\nabla^{2}_{W_{d}}\tilde{\mathcal{L}}_{\text{SDL}}=I_{n_{r}}\otimes(2A).$ (59)
For positive semi-definiteness, we need to show $A\succeq 0$. From the block structure (Claim 3.1) and positive definiteness of $A_{\mathcal{A}\mathcal{A}}$ (Claim 3.2):
$A=\begin{bmatrix}A_{\mathcal{A}\mathcal{A}}&0\\ 0&0\end{bmatrix}\succeq 0,$ (60)
which implies $\nabla^{2}_{W_{d}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})\succeq 0$, proving condition (3).
Step 6: Non-zero loss for polysemantic patterns. We now prove that $\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})>0$ when the pattern exhibits polysemanticity.
By Theorem LABEL:thm:global-necessary-sufficient, $\tilde{\mathcal{L}}_{\text{SDL}}=0$ if and only if:
$w_{r}^{d}=W_{d}\sigma(W_{e}w_{p}^{d})=W_{d}z_{d}\quad\forall d\in[n].$ (61)
Since the pattern has at least one $|\mathcal{F}_{i}|\geq 2$, fix such an $i\in\mathcal{A}$ and choose $d_{1},d_{2}\in\mathcal{F}_{i}$ with $d_{1}\neq d_{2}$. Both features activate neuron $i$, so $(z_{d_{1}})_{i}>0$ and $(z_{d_{2}})_{i}>0$.
The perfect reconstruction conditions for these features are:
$w_{r}^{d_{1}}$ $=\sum_{j=1}^{n_{q}}w_{d}^{j}(z_{d_{1}})_{j}=\sum_{j\in\mathcal{F}_{j}:d_{1}\in\mathcal{F}_{j}}w_{d}^{j}(z_{d_{1}})_{j},$ (62)
$w_{r}^{d_{2}}$ $=\sum_{j=1}^{n_{q}}w_{d}^{j}(z_{d_{2}})_{j}=\sum_{j\in\mathcal{F}_{j}:d_{2}\in\mathcal{F}_{j}}w_{d}^{j}(z_{d_{2}})_{j}.$ (63)
Both sums include the term $w_{d}^{i}(z_{d_{k}})_{i}$ for $k\in\{1,2\}$. However, the coefficients differ: $(z_{d_{1}})_{i}\neq(z_{d_{2}})_{i}$ in general (since different input patterns).
Key observation: Under Assumption 2.4, $w_{r}^{d_{1}}\not\propto w_{r}^{d_{2}}$ for distinct features. If both (61) equations held, we would need:
$\frac{w_{r}^{d_{1}}}{(z_{d_{1}})_{i}}=w_{d}^{i}=\frac{w_{r}^{d_{2}}}{(z_{d_{2}})_{i}},$ (64)
which requires $w_{r}^{d_{1}}\propto w_{r}^{d_{2}}$, contradicting Assumption 2.4.
Therefore, the optimal decoder $W_{d}^{*}$ cannot satisfy (61) for all $d$, giving:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{d}^{*},W_{e})=\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-W_{d}^{*}z_{d}\|^{2}>0,$ (65)
proving condition (1).
Conclusion. We have constructed $(W_{d}^{*},W_{e})$ exhibiting the polysemanticity pattern such that all three conditions hold, completing the proof. ∎
Appendix D Proof of Theorem 3.8
@@
We show that hierarchical concept structures naturally give rise to polysemanticity patterns with feature absorption, and then apply Theorem 3.5 to establish the existence of local minima.
Consider ground-truth features exhibiting hierarchical structure with parent concept $p$ and sub-concepts $\mathcal{C}=\{c_{1},\ldots,c_{k}\}$ satisfying Definition 3.7. Any partition of $\mathcal{C}$ into non-empty subsets $\mathcal{C}_{1},\mathcal{C}_{2}$ with $\mathcal{C}_{1}\cup\mathcal{C}_{2}=\mathcal{C}$ and $\mathcal{C}_{1}\cap\mathcal{C}_{2}=\emptyset$ induces a polysemanticity pattern as follows.
Define $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ where:
  • $\mathcal{F}_{i}=\mathcal{C}_{1}$ for some neuron $i\in[n_{q}]$
  • $\mathcal{F}_{j}=\mathcal{C}_{2}$ for some neuron $j\in[n_{q}]$, $j\neq i$
  • $\mathcal{F}_{\ell}=\emptyset$ for all other neurons $\ell\notin\{i,j\}$
This constitutes a polysemanticity pattern since:
  1. 1.
    $(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ forms a partition of $[n]$ restricted to the sub-concepts (assuming other ground-truth features are assigned to other neurons)
  2. 2.
    At least one of $|\mathcal{F}_{i}|\geq 2$ or $|\mathcal{F}_{j}|\geq 2$ holds, satisfying the polysemanticity condition in Definition 3.4
  3. 3.
    The pattern exhibits feature absorption: neuron $i$ captures a subset of sub-concepts (absorbed feature) while neuron $j$ captures the remaining sub-concepts (main line interpretation)
If this polysemanticity pattern is realizable (Definition 3.4), then by Theorem 3.5, there exists a configuration $(W_{d},W_{e})$ that is a local minimum of $\tilde{\mathcal{L}}_{\text{SDL}}$ exhibiting this pattern. This completes the proof. ∎
Appendix E The Linear Representation Bench
Appendix F Limitations and Future Works
  • Assumption Violations. Our analysis relies on Assumptions 2.3 and 2.4, which may not hold perfectly in realworld neural networks.
  • Extreme Sparsity. Some of our theoretical results are limited to the extreme sparsity ($S\to 1$) assumption, which may not hold in realworld neural networks.
  • Convergence Guarantees. We characterize the global and local minima of the optimization landscape. Future works can provide convergence guarantees or sample complexity bounds following our framework.
  • Classical SDL Methods. Established dictionary learning algorithms have demonstrated success in signal processing but remain unexplored for mechanistic interpretability, representing promising future directions.

### v3

@@ Appendix I Proofs of the Theorems
We prove both directions of the equivalence.
Sufficient condition ($\Rightarrow$). Suppose $w_{r}^{d}=W_{D}\sigma(W_{E}w_{p}^{d})$ for all $d\in[n]$. By definition of the approximate loss:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})$ $=\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\|^{2}$ (66)
$=\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-w_{r}^{d}\|^{2}$ (67)
$=0$ (68)
Necessary condition ($\Leftarrow$). Suppose $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})=0$. Then:
$\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\|^{2}=0$ (69)
Since $M_{d}>0$ for all $d\in[n]$ (by definition, $M_{d}=\Pr(\mathbf{x}(s)=x_{d}(s)e_{d})\cdot\mathbb{E}[x_{d}(s)^{2}|x_{d}(s)>0]$ where both factors are positive under Assumptions 2.3 and 2.4):
$\|w_{r}^{d}-W_{D}\sigma(W_{E}w_{p}^{d})\|^{2}=0\quad\forall d\in[n]$ (70)
Therefore:
$w_{r}^{d}=W_{D}\sigma(W_{E}w_{p}^{d})\quad\forall d\in[n]$ (71)
I.5 Proof of Theorem 3.7
@@
We construct a configuration $(W_{D}^{*},W_{E}^{*})$ exhibiting the polysemanticity pattern and explicitly verify that both gradients vanish.
Step 1: Given encoder $W_{E}^{*}$ from realizability.
Since the activation pattern $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{n_{q}})$ is realizable and forms a partition of $[n]$, by Definition 3.6, there exists an encoder $W_{E}^{*}\in\mathbb{R}^{n_{q}\times n_{p}}$ such that:
$\forall i\in[n_{q}],\quad\mathcal{F}_{i}=\left\{d\in[n]:\left(\sigma(W_{E}^{*}w_{p}^{d})\right)_{i}>0\right\}$ (72)
Fix this $W_{E}^{*}$ and define the activation vectors:
$z_{d}:=\sigma(W_{E}^{*}w_{p}^{d})\in\mathbb{R}^{n_{q}},\quad d=1,\ldots,n$ (73)
Since $\mathcal{P}$ forms a partition, for each feature $d$ there exists a unique neuron $i(d)$ such that $d\in\mathcal{F}_{i(d)}$, and:
$z_{d}=\langle w_{E}^{*,i(d)},w_{p}^{d}\rangle\cdot e_{i(d)}$ (74)
That is, only neuron $i(d)$ activates for feature $d$.
Step 2: Construct optimal decoder $W_{D}^{*}$.
For each neuron $i\in[n_{q}]$, we construct the decoder column $W_{D}^{*,i}$ by minimizing the loss over features activating that neuron.
Dead neurons: If $\mathcal{F}_{i}=\emptyset$, set $W_{D}^{*,i}=0$.
Active neurons: If $\mathcal{F}_{i}\neq\emptyset$, solve:
$W_{D}^{*,i}=\arg\min_{W_{D}^{i}\in\mathbb{R}^{n_{r}}}\sum_{d\in\mathcal{F}_{i}}M_{d}\left\|w_{r}^{d}-W_{D}^{i}(z_{d})_{i}\right\|^{2}$ (75)
This is a least squares problem with solution:
$W_{D}^{*,i}=\frac{\sum_{d\in\mathcal{F}_{i}}M_{d}(z_{d})_{i}w_{r}^{d}}{\sum_{d\in\mathcal{F}_{i}}M_{d}(z_{d})_{i}^{2}}$ (76)
Step 3: Verify $\nabla_{W_{D}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=0$.
The loss is $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D},W_{E})=\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-W_{D}z_{d}\|^{2}$. The gradient with respect to column $W_{D}^{i}$ is:
$\nabla_{W_{D}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}=-2\sum_{d=1}^{n}M_{d}(w_{r}^{d}-W_{D}z_{d})(z_{d})_{i}$ (77)
Since $\mathcal{P}$ is a partition, $(z_{d})_{i}\neq 0$ only when $i=i(d)$, i.e., when $d\in\mathcal{F}_{i}$. Therefore:
$\nabla_{W_{D}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=-2\sum_{d\in\mathcal{F}_{i}}M_{d}(w_{r}^{d}-W_{D}^{*,i}(z_{d})_{i})(z_{d})_{i}$ (78)
By construction (76), $W_{D}^{*,i}$ satisfies the normal equation:
$\sum_{d\in\mathcal{F}_{i}}M_{d}(z_{d})_{i}(w_{r}^{d}-W_{D}^{*,i}(z_{d})_{i})=0$ (79)
Therefore $\nabla_{W_{D}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=0$ for all $i\in[n_{q}]$.
Step 4: Verify $\nabla_{W_{E}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=0$.
Within the activation pattern region $\Omega_{\mathcal{A}}$, for feature $d\in\mathcal{F}_{i}$, the activation is:
$z_{d}=\langle w_{E}^{i},w_{p}^{d}\rangle\cdot e_{i}$ (80)
The gradient with respect to encoder row $w_{E}^{i}$ is:
$\nabla_{w_{E}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}=\sum_{d=1}^{n}M_{d}\nabla_{w_{E}^{i}}\left\|w_{r}^{d}-W_{D}z_{d}\right\|^{2}$ (81)
Only features $d\in\mathcal{F}_{i}$ have non-zero contribution. For such $d$:
$\frac{\partial z_{d}}{\partial w_{E}^{i}}=w_{p}^{d}\otimes e_{i}$ (82)
where $\otimes$ denotes outer product. Therefore:
$\nabla_{w_{E}^{i}}\left\|w_{r}^{d}-W_{D}z_{d}\right\|^{2}$ $=-2(w_{r}^{d}-W_{D}z_{d})^{\top}W_{D}\frac{\partial z_{d}}{\partial w_{E}^{i}}$ (83)
$=-2(w_{r}^{d}-W_{D}^{*,i}\langle w_{E}^{*,i},w_{p}^{d}\rangle)^{\top}W_{D}^{*,i}w_{p}^{d}$ (84)
The total gradient is:
$\nabla_{w_{E}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}$ $=-2\sum_{d\in\mathcal{F}_{i}}M_{d}(w_{r}^{d}-W_{D}^{*,i}\langle w_{E}^{*,i},w_{p}^{d}\rangle)^{\top}W_{D}^{*,i}w_{p}^{d}$ (85)
$=-2(W_{D}^{*,i})^{\top}\sum_{d\in\mathcal{F}_{i}}M_{d}\langle w_{E}^{*,i},w_{p}^{d}\rangle(w_{r}^{d}-W_{D}^{*,i}\langle w_{E}^{*,i},w_{p}^{d}\rangle)$ (86)
From the normal equation in Step 3 with $(z_{d})_{i}=\langle w_{E}^{*,i},w_{p}^{d}\rangle$:
$\sum_{d\in\mathcal{F}_{i}}M_{d}\langle w_{E}^{*,i},w_{p}^{d}\rangle(w_{r}^{d}-W_{D}^{*,i}\langle w_{E}^{*,i},w_{p}^{d}\rangle)=0$ (87)
Therefore:
$\nabla_{w_{E}^{i}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=-2(W_{D}^{*,i})^{\top}\cdot 0=0$ (88)
This holds for all active neurons. For dead neurons ($\mathcal{F}_{i}=\emptyset$), the gradient is automatically zero.
Step 5: $(W_{D}^{*},W_{E}^{*})$ is a partial optimum.
By Theorem 3.2, $\tilde{\mathcal{L}}_{\text{SDL}}$ is convex in $W_{D}$ for fixed $W_{E}^{*}$, and convex in $W_{E}$ for fixed $W_{D}^{*}$ within $\Omega_{\mathcal{A}}$.
Since we have verified:
  • $\nabla_{W_{D}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=0$ (Step 3)
  • $\nabla_{W_{E}}\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=0$ within $\Omega_{\mathcal{A}}$ (Step 4)
By the first-order optimality condition for convex functions:
  • $W_{D}^{*}$ minimizes $\tilde{\mathcal{L}}_{\text{SDL}}(\cdot,W_{E}^{*})$
  • $W_{E}^{*}$ minimizes $\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},\cdot)$ over $\Omega_{\mathcal{A}}$
By standard results in biconvex optimization (Gorski et al., 2007), $(W_{D}^{*},W_{E}^{*})$ is a partial optimum.
Step 6: Non-zero loss.
Since the pattern is polysemantic, there exists $i\in[n_{q}]$ with $|\mathcal{F}_{i}|\geq 2$. Choose distinct $d_{1},d_{2}\in\mathcal{F}_{i}$.
By Theorem 3.4, zero loss requires:
$w_{r}^{d_{1}}=W_{D}^{*}z_{d_{1}}=W_{D}^{*,i}(z_{d_{1}})_{i},\quad w_{r}^{d_{2}}=W_{D}^{*}z_{d_{2}}=W_{D}^{*,i}(z_{d_{2}})_{i}$ (89)
If both equations held, then:
$\frac{w_{r}^{d_{1}}}{(z_{d_{1}})_{i}}=W_{D}^{*,i}=\frac{w_{r}^{d_{2}}}{(z_{d_{2}})_{i}}$ (90)
Since $(z_{d_{1}})_{i}=\langle w_{E}^{*,i},w_{p}^{d_{1}}\rangle>0$ and $(z_{d_{2}})_{i}=\langle w_{E}^{*,i},w_{p}^{d_{2}}\rangle>0$, this requires:
$w_{r}^{d_{1}}=\frac{(z_{d_{1}})_{i}}{(z_{d_{2}})_{i}}w_{r}^{d_{2}}$ (91)
This means $w_{r}^{d_{1}}\propto w_{r}^{d_{2}}$, contradicting Assumption 2.4 (distinct features have linearly independent reconstruction vectors). Therefore:
$\tilde{\mathcal{L}}_{\text{SDL}}(W_{D}^{*},W_{E}^{*})=\sum_{d=1}^{n}M_{d}\|w_{r}^{d}-W_{D}^{*}z_{d}\|^{2}>0$ (92)
Conclusion. We have constructed $(W_{D}^{*},W_{E}^{*})$ exhibiting the polysemanticity pattern, explicitly verified both gradients vanish, and shown it achieves positive loss, completing the proof. ∎
I.6 Proof of Theorem 3.10
Proof.
We construct a realizable pattern exhibiting feature absorption by scaling down the parent neuron’s encoder until one sub-concept separates, then adding a dedicated neuron for the separated feature.
Step 1: Initial configuration and feature geometry.
Since $\mathcal{P}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{M})$ is realizable, there exists an encoder $W_{E}\in\mathbb{R}^{M\times n_{p}}$ and threshold $c>0$ such that:
$\forall i\in[M],\quad\mathcal{F}_{i}=\{d\in[n]:\langle w_{E}^{i},w_{p}^{d}\rangle>c\}$ (93)
Fix any parent neuron $i^{*}\in[M]$ with sub-concepts $\mathcal{F}_{i^{*}}=\{d_{i^{*},1},\ldots,d_{i^{*},k_{i^{*}}}\}$ where $k_{i^{*}}\geq 2$.
For each sub-concept, compute its activation strength:
$a_{j}:=\langle w_{E}^{i^{*}},w_{p}^{d_{i^{*},j}}\rangle>c,\quad\forall j\in[k_{i^{*}}]$ (94)
Since there are finitely many sub-concepts, define:
$a_{\min}:=\min_{j\in[k_{i^{*}}]}a_{j}>c$ (95)
Let $j^{*}\in[k_{i^{*}}]$ be the index achieving this minimum: $a_{j^{*}}=a_{\min}$.
Step 2: Scale down parent neuron to separate one sub-concept.
We shrink the encoder row $w_{E}^{i^{*}}$ by multiplying with scaling factor $\lambda\in(0,1)$:
$\tilde{w}_{E}^{i^{*}}:=\lambda\cdot w_{E}^{i^{*}}$ (96)
The new activation strengths become:
$\tilde{a}_{j}:=\langle\tilde{w}_{E}^{i^{*}},w_{p}^{d_{i^{*},j}}\rangle=\lambda\cdot a_{j}$ (97)
Choose $\lambda$ such that:
$\lambda:=\frac{c+a_{\min}}{2a_{\min}}\in\left(\frac{c}{a_{\min}},1\right)$ (98)
This gives:
  • For the minimum-activation feature: $\tilde{a}_{j^{*}}=\lambda a_{\min}=\frac{c+a_{\min}}{2}<c+\epsilon$ for any small $\epsilon>0$
  • For all other features: $\tilde{a}_{j}=\lambda a_{j}>\lambda a_{\min}=\frac{c+a_{\min}}{2}$
By choosing the threshold to be $c^{\prime}=\frac{c+a_{\min}}{2}$, we have:
[…]
Slightly increasing $c^{\prime}$ by infinitesimal $\epsilon>0$ gives $c^{\prime\prime}=c^{\prime}+\epsilon$, ensuring:
$\tilde{a}_{j^{*}}<c^{\prime\prime}<\tilde{a}_{j}\quad\forall j\neq j^{*}$ (100)
Therefore, with encoder row $\tilde{w}_{E}^{i^{*}}$ and threshold $c^{\prime\prime}$:
$\{d\in[n]:\langle\tilde{w}_{E}^{i^{*}},w_{p}^{d}\rangle>c^{\prime\prime}\}=\mathcal{F}_{i^{*}}\setminus\{d_{i^{*},j^{*}}\}$ (101)
Step 3: Construct dedicated neuron for separated feature.
We create a new encoder row parallel to the separated feature:
$w_{E}^{new}:=\alpha\cdot w_{p}^{d_{i^{*},j^{*}}}$ (102)
where $\alpha>0$ is chosen sufficiently large.
By the unit-norm assumption ($\|w_{p}^{d_{i^{*},j^{*}}}\|=1$):
$\langle w_{E}^{new},w_{p}^{d_{i^{*},j^{*}}}\rangle=\alpha\|w_{p}^{d_{i^{*},j^{*}}}\|^{2}=\alpha$ (103)
For any other feature $d\neq d_{i^{*},j^{*}}$:
$\langle w_{E}^{new},w_{p}^{d}\rangle=\alpha\langle w_{p}^{d_{i^{*},j^{*}}},w_{p}^{d}\rangle\leq\alpha M$ (104)
where $M<1$ is the maximum interference from Assumption 2.4.
Choose $\alpha$ large enough such that:
$\alpha>c^{\prime\prime},\quad\alpha M<c^{\prime\prime}$ (105)
This is possible by taking $\alpha=\frac{2c^{\prime\prime}}{1+M}$, giving:
$\alpha=\frac{2c^{\prime\prime}}{1+M}>c^{\prime\prime},\quad\alpha M=\frac{2c^{\prime\prime}M}{1+M}<c^{\prime\prime}$ (106)
where the first inequality uses $M<1$ and the second uses $M<1$.
Therefore:
  • $\langle w_{E}^{new},w_{p}^{d_{i^{*},j^{*}}}\rangle=\alpha>c^{\prime\prime}$ (separated feature activates new neuron)
  • $\langle w_{E}^{new},w_{p}^{d}\rangle\leq\alpha M<c^{\prime\prime}$ for all $d\neq d_{i^{*},j^{*}}$ (no other features activate)
Step 4: The new activation pattern.
Construct the expanded encoder $W_{E}^{\prime}\in\mathbb{R}^{(M+1)\times n_{p}}$:
$W_{E}^{\prime}=\begin{bmatrix}w_{E}^{1}\\ \vdots\\ w_{E}^{i^{*}-1}\\ \tilde{w}_{E}^{i^{*}}\\ w_{E}^{i^{*}+1}\\ \vdots\\ w_{E}^{M}\\ w_{E}^{new}\end{bmatrix}$ (107)
With threshold $c^{\prime\prime}$, the resulting activation pattern is:
$\mathcal{P}^{\prime}=(\mathcal{F}_{1},\ldots,\mathcal{F}_{i^{*}-1},\mathcal{F}_{i^{*}}\setminus\{d_{i^{*},j^{*}}\},\mathcal{F}_{i^{*}+1},\ldots,\mathcal{F}_{M},\{d_{i^{*},j^{*}}\})$ (108)
This pattern is realizable by construction. ∎